OpenStudy (anonymous):

Find the horizontal and vertical asymptotes of the curve. Y= x^2 + 3/3x^2 − 26x − 9

OpenStudy (anonymous):

Hey loco!! Where have you been man?!

OpenStudy (anonymous):

Anwar!!! What's up? Busy man.

OpenStudy (anonymous):

Still have one more final exam to do, then I can go see my family for the first time in 17 weeks.

OpenStudy (anonymous):

Oh... wow. I'm taking Calc I as a summer class.

OpenStudy (anonymous):

Oh your summer session has already started?!

OpenStudy (anonymous):

Yeah, going into the 3rd week.

OpenStudy (anonymous):

Let's get to your question now. Is the function \(y=\frac{x^2 + 3}{3x^2 − 26x − 9}\)?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Good. So let's start by vertical asymptotes. The vertical asymptotes of a rational function occurs at the zeros of the denominator. So, we have to find when the denominator is \(0\) by solving \(3x2−26x−9=0 \implies (3x+1)(x-9)=0\) That's \(x=-\frac{1}{3}\) or \(x=9\). Since non of them is a zero of the numirator then these two values would be your vertical asymptotes.

OpenStudy (anonymous):

For horizontal asymptotes, you have to take the limit of the function as it goes to infinity, that's: \[\lim_{x \rightarrow \infty}{x^2+3 \over 3x^2-26x-9}\], the limit of this is just the coffieceint of the hieghst degree of x in the numerator over the coffieceint of the hieghst degree in the denominator, which is \(\frac{1}{3}\).

OpenStudy (anonymous):

I don't understand the horizontal asymptote part.

OpenStudy (anonymous):

When you want to find the horizontal asymptotes of a function, you're actually studying its behavior as it goes to infinty. So we took the limit as x goes to infinity. Now, let's focus on how we can find the limit of a rational function, as ours, when it's approaching infinty. In our case the degree of the numerator is the same as the degree of denominator. So, the limit will be just the coefficients of x^2, which are 1 in the numerator and 3 in the denominator.

OpenStudy (anonymous):

Oh ok... good explanation.

OpenStudy (anonymous):

Really?! :D

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Find the limit. (If you need to use or –, enter INFINITY or –INFINITY, respectively.) http://www.webassign.net/cgi-bin/symimage.cgi?expr=lim_%28x-%3Einfinity%29%20text%28%20%20%29%20%28x%2B2%29%2Fsqrt%284%20x%5E2%2B1%29

OpenStudy (anonymous):

Can you help me with this one?

OpenStudy (anonymous):

I have to go now! I'll help you when I come back.

OpenStudy (anonymous):

I got it. I have another problem though. Thanks!